For the function, f(x) = x2 −2 defined over [1,5], verify : L(P, f ) ≤ U(−P, f )

where P is the partition which divides [1,5] into four equal intervals.


1
Expert's answer
2021-06-15T06:31:04-0400

When P is the position that divides [1,5] into four equal parts. The position of [1,5] into these parts are, "P=[I_1, I_2, I_3, I_4]"

Where I1=[1,2], I2=[2,3],I3=[3,4],I4=[4,5]

Now the length of each interval is given by

"S_1=2-1=1= \triangle x_1"

"S_2=3-2=1= \triangle x_2"

"S_3=4-3=1= \triangle x_3"

"S_4=5-4=1= \triangle x_4"

Now the downbound of the function f(x) corresponding to each interval is given by

"\therefore f(x)=x^2-2"

"m_1=1^2-2=-1"

"m_2=2^2-2=2"

"m_3=3^2-2=7"

"m_4=4^2-2=14"

"L(p,f)= \sum_{r=1}^4 m_r \triangle x_r =m_1 \triangle x_1+m_2 \triangle x_2+m_3 \triangle x_3+m_4\triangle x_4"

"L(p,f) =-1*1+2*1+7*1+14*1 =-1+2+7+14=22"


Now "-P=[-I_1, -I_2, -I_3, -I_4]"

Where I'1=[-1,-2], I'2=[-2,-3],I'3=[-3,-4],I'4=[-4,-5]

The length of each interval = "1 = \triangle x_r \space \space \space 1 \eqslantless r \eqslantless 4"

The upperbound of "f(x)=x^2-2" over each interval is given by

"\therefore f(x)=x^2-2"

"m_1=(-1)^2-2=-1"

"m_2=(-2)^2-2=2"

"m_3=(-3)^2-2=7"

"m_4=(-4)^2-2=14"

The upper sum is defined as

"U(-p,f)= \sum_{r=1}^4 m_r \triangle x_r"

"U(-p,f) =-1*1+2*1+7*1+14*1 =-1+2+7+14=22"

Hence "L(p,f)=U(-p,f)" verified

We also know that the upper sum is always greater than the lower sum, hence

"L(p,f)\eqslantless U(-p,f)" verified